In a mass spectrum, two neighboring ion peaks (both ions singly charged) were at a

distance of 0.80 m/z units. The first peak had a mass-to-charge ratio of 480. From this information it

follows that the MS instrument has this value for the resolving power.( in detail please)
Answer #1

Resolving power = (m/z) / \DeltaM

= 480 / 0.80 = 600